# A Survey on Computational Solutions for Reconstructing Complete Objects by Reassembling Their Fractured Parts

### v1

@@ 4.3 Joint Segmentation and Matching
Figure 10: Joint matching and segmentation [61].
@@
In the context of fractured object reassembly, the only approach that explicitly models this objective is JIGSAW [61] (see Figure 10), It extracts point-wise features using transformers, which are used to predict point labels of original or fracture surfaces and perform point-wise matching. Joint segmentation and matching are achieved using shared deep point-wise feature descriptors.
5 Template Shape Space
@@
Early approaches (see Figure 11(a)) include Morphable faces [166] and SCAPE [68], which focus on 3D shape spaces with the same topology. Recent examples along this line include SMPL [155] for humans, SMAL [167] for animals, MANO [168] of hands, and 3DMM [169] for faces, which are learned from larger datasets and therefore exhibit significantly improved performance in solving inverse problems. These models decouple shape variations and pose variations. Shape variations are derived from PCA analysis. Pose variations, e.g., in SMAL and SMPL are encoded using bones. A limitation of these approaches is that require consist meshes to learn the shape prior.
@@
DeepSDF [157] and follow-up work show that shape generative models can be learned from raw data points with correspondences. SALD [175] shows that MLP-based implicit surfaces can also encode deformable shapes, but generalization of shapes that undergo large pose changes is very limited. GenCorres [139] addresses the generalization issue under large pose changes by developing an approach to compute dense correspondences between adjacent implicit surfaces, allowing us to apply ARAPReg to define regularization losses. It leads to significantly improved synthetic shapes under novel poses.
@@
One way to address this issue is to learn generative under explicit representations such as point clouds and meshes. Many approaches have shown promising results for point cloud generators for man-made shapes [193, 194, 195, 160, 196]. Although it is possible to fit a point cloud generator to shape collections with large geometric and topological variations, there are two limitations that prevent them from serving template models for fractured object reassembly. First, most approaches require the shapes to be consistently oriented, which only applies to a very limited number of categories. Generalization behavior under unaligned training instances is much worse. Second, the correspondences induced from the point cloud generator are not good enough across the synthetic shapes, making it difficult to properly define correspondences between the dynamic template and the input fragments. Recent results in mesh representation, such as deep marching tetrahedra [197], MeshGPT [198], and MeshAnything [199], also show impressive synthesis results. However, the two limitations of the point cloud representation remain under the mesh representation. An open question is how to address these two limitations.
6 Template-Based Reassembly
6.1 Complete object priors
In addition to feature-based reassembly, a complete object model from the same or a similar class can provide valuable prior knowledge to facilitate reassembly. For example, when reassembling objects such as skulls, pottery, and architectural elements, complete models are often available and can serve as templates to infer the initial arrangement of fragments. In recent data-driven techniques, these global geometric priors can also be learned from collected data sets using neural networks.
@@ 6.1 Complete object priors
Methods using local-global alignment match fragmented pieces $x_{i}$ to the global prior $P$ and use this to obtain the initial or rough arrangement of the fragments. Their effectiveness depends on three factors: (1) the design of $P$, (2) the modeling/extraction of features from $\{x_{i}\}$ and $P$ for their matching, and (3) the alignment. The global prior $P$ could be a selected/computed explicit template or learned implicit parameters. Classical geometric reassembly methods often search for a good explicit shape and use it as a template or global prior, then design partial matching schemes to find local-global alignments. In contrast, as discussed in Sec. 6.3, recent learning-based methods tend to train neural networks to build implicit shape priors from datasets, and then train/optimize the encoding of both local fragments and global template so that their latent codes can infer the matching and alignment of the fragments.
@@
In classical geometric methods, to support effective local-global alignment, the explicit shape template is often selected as an average or typical shape, or a best-matched object retrieved from a database. Using a typical object as prior shape is common in skull and cultural heritage reassembly tasks. For example, [200, 201, 189] used a standard human skull as $P$ to provide alignment guidance. [202] used a complete model from the museum’s collection as the prior to define the 3D support surface of the fragmented disk. When a larger collection of complete exemplar objects is available, retrieving a template object with the most similar geometry can provide a more adaptive global prior. For example, in the pottery reassembly field, researchers have built many digital databases of pottery. They support various types of local-global matching for the given fragments, such as drawing sketch [203], a portion of the object geometry [204], and curve descriptors [87, 205] to retrieve relevant potteries as priors $P$. In the archaeological and architectural reconstruction of a ruined monument task, to construct the most similar geometry as prior $P$, [50, 206] conducted a careful study of historical texts that mention the building under consideration. Given the selected shape prior $P$, as for the second factor, feature extraction computes the keypoints and their descriptors on both fragments and the global prior, such as depth and axial symmetry descriptor [203], salient points [204], parabolic contour [87], spin images [200, 201], Signature of Histograms of OrienTations (SHOT) descriptor [189]. With the extracted keypoints and features, fragment transformations are computed to roughly assemble the input parts to the template. Researchers developed various alignment methods using keypoint correspondences, such as iterative closed points (ICP) [200], geometric distance [201], local affine-invariant 3D moments [87], Earth mover’s distance [204], and RANSAC algorithm [189].
@@
Implicit prior for local-global alignment is learned from an annotated 3D dataset to solve the partial matching from the fragments to the global shape. The shape prior is encoded in a neural network, which is trained on a 3D segmentation dataset, that the weights are adjusted to learn the semantic relationship between the fragments and the complete objects. In the inference stage, given a set of parts $X$, trained networks are used as encoders to encode the parts into latent codes, which inherit the implicit shape prior to the possible complete object $P$. Chaudhuri et al. [207] trained a Bayesian network on a 3D segmentation dataset to encode semantic and geometric relationships among shape components. Then a probabilistic model is used to compose and align the input parts. Mo et al. [190] proposed a more comprehensive 3D object dataset, named PartNet, a large-scale dataset of 3D objects annotated with hierarchical and instance-level 3D part information. Following the PartNet dataset, Zhan et al. [208] and Narayan et al. [209] trained an encoder on it to encode the query parts to semantic latent codes, which inherit the shape prior to the complete object. They framed the 3D part assembly problem as a graph-learning problem. They took the latent codes as input and used graph neural networks to assemble the parts into complete objects. They demonstrated their work on 3D chairs, tables, and lamps reassembly tasks.
@@
Global evaluation as the complete object prior $P$ provides a regularization term to constrain the reassembled results. It has two variants: hand-crafted functions or learnable discriminators. Hand-crafted metrics are widely adopted in earlier reassembly tasks. In the image stitching task, an entropy-based metric is proposed to assess the quality of the stitched panoramic image [210]. In the document reassembly task, Optical Character Recognition (OCR) for word or sentence detection [211, 212]. These methods use pair-wise features to suggest candidate alignments of the fragments. They keep the candidate alignments when the evaluation results are promising, while they discard the alignments when the evaluation results are poor. Recent approaches used the capacity of large-scale datasets, trained neural networks as discriminators, and used them for global evaluation. The discriminator distinguishes the reassembled the results and the natural objects to push the network propose a better reassemble configuration that makes the results look natural. Li et al. [191] propose a GAN to solve the Jigsaw puzzles. The designed a network-based discriminator and an adversarial loss to classify the an image is reassembled or real. The adversarial loss pushes the GAN to assemble images from pieces that looks natural. Then a discriminator was trained as adversarial loss to optimize the mating configuration that should look visually realistic like an object.
@@
In summary, local-global alignment often works well on a specific category that has a typical complete object or a large enough database as reference. The quality of the initial alignment depends on how to encode the part-global relationship into features. Due to the lack of well-annotated database, local-global alignment might be hard to generalize to other object reassembly tasks. In contrast, the global evaluation function provides a global shape guidance to regularize the reassembled results. However, finding a proper evaluation function for efficient reassembly is still challenging.
@@ 6.2.1 Symmetry Detection
Axial symmetry and reflectional symmetry are the two commonly studied symmetry types in geometry analysis and processing, and especially in 3D reassembly.
As shown in Figure 13 (top), axially symmetric, also referred to as rotationally symmetric, indicates that if the complete object’s exterior surface is intersected by a plane perpendicular to its axis, the resulting cross section is a circle or closely resembles one. As a prominent geometrical property, rotational symmetry can serve as a feature to guide the 3D reassembly of fragmented relics, such as sherds. Numerous studies explored the estimation of the axis of symmetry. Earlier studies estimate the rotational axis using differential geometric properties, such as the sphere of curvature [216] and normal intersection [217]. Subsequently, enhancements of the estimated rotational axis have been proposed using M-estimators [218] or Bayesian approaches [219, 220]. However, these methods can be sensitive to geometric deformation, incompleteness, or noise, often demonstrated in fragmented pieces, leading to inaccurate detection of an object’s symmetrical axis. [221] suggested a workflow for axis estimation based on minimum thickness direction. This method does not rely on differential geometric properties, making it resilient to local defectiveness, noise, and outliers. Recent studies [222, 223, 213] have broadly adopted the random sample consensus algorithm (RANSAC) to find good initial estimates. [222] extended the Cao and Mumford’s method [216], to consider both the inner and outer surfaces. This technique introduces PotSAC, a two-stage axis estimator based on a variant of the RANSAC algorithm followed by a robust nonlinear least-square refinement, making it capable of handling noisy surface normals. [223] applied beam search to explore multitudes of registration possibilities, addressing false matches between sherds inflicted by indistinctive sharp fracture surfaces. [213] showed that the identification of the symmetry axis from surface normal lines may yield two meaningful local minima. A multiple-solutions RANSAC algorithm is proposed for initial estimates, followed by a coordinate-descent algorithm to refine these estimates. The authors showed that their method ensures convergence, achieves a faster convergence rate, and provides state-of-the-art performance in axis estimation.
As illustrated in Figure 13 (bottom), reflectional symmetry refers to the property that an object does not change upon undergoing a reflection. Human skull typically exhibits reflectional symmetry along the mid-sagittal plane [224]. This kind of symmetry is one of the main geometric priors used in skull reassembly and completion [201, 200].
@@ 6.2.2 Symmetry-guided reassembly
With an estimated symmetry axis, the consistency of the symmetry axis and the axis profile curve can be viewed as a constraint or optimization objective during reassembling. [225, 226] implemented a global constraint that all fragments share the same axis, greatly reducing computational costs and improving noise resistance. This method reduces the search space from 6 DOF (for two fragments) to 1 DOF (relative position of two pieces along the axis of symmetry). [223] relaxed the above constraint by encoding the information about the axis as a descriptor into each break line for efficient reassembly.
@@ 6.3 Learned Shape Spaces
Compared to approaches that rely on similar shapes or symmetric templates to guide fractured object reassembly, few methods have explored the use of a generic template shape space learned from the data to enhance this process. One notable exception is TAssembly [214], which learns a linear shape space specifically for bones. It formulates fractured object reassembly as an optimization problem, adjusting fragment poses and the underlying complete shape in the learned shape space to minimize the distance between the fragments and the dynamic template. However, this method is limited to bone shape spaces, where consistent mesh parameterization makes it relatively straightforward to learn descriptors for predicting dense correspondences. Another promising approach is Jigsaw++ [215], which employs a point cloud generation model with a rectified flow-based retargeting algorithm to infer the complete shape from a partially assembled input. Although this approach shows significant potential for reconstructing daily objects, scaling it to handle more diverse or even open-vocabulary input shapes requires larger models and more data. The challenge remains to generate shape priors when the underlying shape space exhibits large geometric and topological variations. Another limitation is that it does not compute correspondences between the reconstruction and the input fragments. Therefore, it does not enforce the rigidity constraints among fragments.
@@ 7 Applications
We discuss five major application areas of fractured object reassembly (Section 7.1 to Section 7.5) and the relevant application in shape modeling via part assembly.
@@ 7.1 Bone Reduction
3D bone reduction serves as a key step in preoperative surgical planning [227, 228, 229] and constitutes a critical element of intraoperative navigation to reduce fractures [230, 231] or a robotic system [232, 233](see Figure 15). The objective of bone reduction is to restore the native anatomical configuration of bony structures, encompassing both individual bones and articulations. Unlike other fractured entities, fractured bones exhibit the following characteristics. Primarily, human bones exhibit pronounced shape priors, which can be leveraged to facilitate the reassembly procedure (see Section 6). Owing to the inherent reflection symmetry of human anatomy, the intact contralateral bones can serve as the archetypal reference for the pre-fracture bone structure. Secondly, unlike other objects, bone is composed of compact (cortical) bone, cancellous tissue, and cartilage. In instances where the fracture occurs within the cancellous tissue, especially within the metaphyseal region of a long bone, the resultant fracture interface commonly presents considerable indeterminacy, which affects the precise alignment of bone fragments [214, 234]. In the literature, the general workflow of bone reduction includes four components, i.e., CT bone segmentation, template matching, pairwise fragment matching, and multi-fragment matching.
@@
CT bone segmentation endeavors to delineate the target fractured bone from adjacent healthy bone and to discriminate the bone fragments from each other in CT images. Subsequent to delineation, bone fragments are commonly divided into the fracture zone and the intact zone, a process that constitutes one of the most arduous challenges in CT bone segmentation. The methods identifying fracture zone are comprised of automatic approaches such as watershed [227], probabilistic watershed [237], template-matching [234] or semi-automatic methods such as region growing [238, 239] or continuous max-flow [230]. In addition to the non-DL methods mentioned above,  [240] implements supervised contrastive learning with a novel Siamese deep neural network so as to identify and localize the pelvic fracture zones. In [234], an approach was introduced to synergistically segment and reassemble the fragments utilizing the contralateral bone as a template and exploiting the synergistic interaction between segmentation and reduction (see Figure 16). In recent advancements delineated in [241], an innovative dual-stream learning framework has been introduced for the automated segmentation and categorical labeling of pelvic fractures. This framework facilitates the identification of pelvic fracture fragments in varied quantities and anatomical locations through a dual-branch architecture that harnesses distance learning based on bone fragments. Considering that the bone template is constituted by the outer surface of the intact zone, aligning the bone fragment with the template necessitates the segmentation of the outer surface of the intact zone [237, 234]. The bone surface can be extracted by the MarchingCube algorithm [47].
Template matching is employed as an initial coarse reduction step by aligning the bone fragments with the bone template. The healthy contralateral mirrored bone frequently serves as a reference template for the reduction of the fractured side [227, 234, 242, 237, 243, 244]. Matching the bone fragments with the contralateral bone template can be formulated as a part-to-whole matching problem for 3D rigid objects. However, the existence of intra-individual variations, particularly discrepancies in axial torsion, led [245] to conclude that preoperative planning methodologies aimed at reconstructing intricate proximal humerus fractures should not depend uncritically on contralateral anatomical structures. Furthermore, contralateral bone is not always available due to ethical considerations in medical practice and instances where the pathology affects both sides. Due to the similar morphological characteristics of human bones, the parametric statistical shape model (SSM) serves as an alternative template for fracture reduction [214, 228, 230]. The alignment of bone fragments with the parametric morphable template is formulated as a part-to-whole problem for 3D non-rigid objects, whereby the output encompasses the rigid 6D pose of each bone fragment and the deformation parameters of the SSM model. Moreover, to address the complexities associated with joint dislocation, a statistical pose model (SPM) has been meticulously developed to accurately represent the anatomical pose of the bones comprising the bony joint [246]. Significantly, the seminal work by [230] integrates both the SSM and the SPM frameworks to address intra-bone fractures and inter-bone dislocations. Given the inherently non-convex nature of this complex problem, the iterative procedure alternates through three critical stages: fragment alignment update, SSM adaptation, and SPM adaptation.
@@
Pair-wise matching aims at aligning two fragments so that their fracture boundaries can perfectly match. Pair-wise matching can be divided into two steps: identification of fracture boundary and matching of the fracture boundary. After template matching, the fracture boundaries between fragments are close. Subsequently, the Euclidean distance between the surface of two fragments can be used as a criterion to identify the fracture boundary [227]. In instances where fractures occur within rigid regions such as cortical bones, the resulting fracture boundary typically manifests pronounced curvature. Subsequent studies [242, 247, 248, 249, 237, 250] have used this characteristic to effectively identify fracture boundaries. In contrast, in cases where fractures occur within more flexible regions, such as trabecular bone, the fracture surface tends to appear relatively planar, making the fracture boundary less distinct. In such circumstances, interactive annotation of the fracture boundary becomes indispensable [251, 242]. In the context of matching fracture boundaries, if the fracture occurs within the rigid cortical bone, the methodologies for pair-wise matching bear significant resemblance to those employed in the reassembly of precise artifacts, such as cultural relics. These techniques [83, 238, 249] frequently use criteria in which tangential contact points on the respective bone fragments converge at the coordinates of the coincident point, with surface normals oriented oppositely, and matching surface curvatures exhibiting inverse signs. Concurrently, the Iterative Closest Point (ICP) algorithm is an effective instrument to execute pairwise matching [252, 242, 248, 250, 227]. In contrast, in instances where the fracture occurs within the more pliable, trabecular bone, or in scenarios where diminutive fragments are absent near the boundary, pair-wise matching is likely to be ineffective. In such circumstances, the previously discussed template matching emerges as an essential and indispensable methodology for bone reduction. Note that while common pairwise registrations and pairwise matching techniques such as ICP and feature matching are used in pairwise bone fragment matching, the approaches introduced above instill domain-specific knowledge of individual bones.
Multi-fragment matching, recast as a quintessential NP-hard problem, seeks to globally optimize the poses of multiple fragments. The input for multi-fragment matching includes the current poses of all bone fragments following template matching and pair-wise matching. The objective of global optimization is to maximize the region matching score, which combines the scores of template matching (if a template exists) and pair-wise matching [189]. To achieve this, existing methodologies typically employ an exhaustive search [238] or a greedy search strategy [253] incorporating heuristic rules to address the aforementioned global optimization. Within the greedy search strategy, bone fragments from the inactive set are sequentially added to the expanding set of already aligned objects, while ensuring the correct alignment of objects within the active set. The next fragment in the inactive set is selected based on the highest score, subject to penetration validity checks. In [189], the score was calculated as the weighted sum of the scores from prior template matching and pair-wise matching. Furthermore, metrics such as the distance between the fragment and the active set [247], the area of the fracture surface [250], and the size of the bone fragment [227] were also used as priority scoring criteria. To avoid entrapment in local minima, multiple fragments can be chosen, rather than a single fragment, to add to the active set using the beam search approach [189]. A notable challenge in multi-fragment matching is the potential accumulation of pairwise/template errors in multi-piece alignment. To address this issue,  [247, 237] introduced the ideal mates concept, which enables the calculation of the multi-piece alignment error without depending on the pairwise registration error [253]. Unlike [247, 237][189] implemented graph optimization on the target graph, representing the final correspondences among the assembled bone fragments, to globally refine existing transformations. Similar to pairwise matching, there are similarities between the multi-fragment matching approaches introduced for bone reduction and those used in restoring stone fragments. However, most approaches to bone reduction consider domain-specific knowledge.

### v2

@@ 4.3 Joint Segmentation and Matching
Figure 11: Joint segmentation and matching [62].
In the context of fractured object reassembly, JIGSAW [62] (see Figure 11) pioneered a deep learning approach for joint segmentation and matching. It effectively utilizes transformer architecture to extract point-wise features and proposed primal-dual descriptors for complementary matching. JIGSAW also demonstrates how learning approaches can be effectively combined with classical geometric algorithms. By integrating learned point-wise features with RANSAC and transformation synchronization methods, it achieves robust pose estimation. This hybrid approach leverages both the pattern recognition capabilities of deep learning and the geometric guarantees of classical methods, suggesting that combining complementary techniques may lead to more effective solutions.
5 Template Shape Space
@@
Early approaches (see Figure 12(a)) include Morphable faces [170] and SCAPE [69], which focus on 3D shape spaces with the same topology. Recent examples along this line include SMPL [164] for humans, SMAL [171] for animals, MANO [172] of hands, and 3DMM [173] for faces. These models, trained on larger datasets, demonstrate significantly improved performance in solving inverse problems. They decompose shape variations from pose variations - shape variations are captured through PCA analysis, while pose variations (as in SMAL and SMPL) are encoded using skeletal structures. However, a key limitation of these approaches is that they require training data with consistent mesh topology, meaning that all meshes must share the same vertex structure.
@@
DeepSDF [166] and follow-up work show that shape generative models can be learned from raw data points without correspondences. SALD [179] shows that MLP-based implicit surfaces can also encode deformable shapes, but generalization of shapes that undergo large pose changes is very limited. GenCorres [143] addresses the generalization issue under large pose changes by developing an approach to compute dense correspondences between adjacent implicit surfaces, allowing us to apply ARAPReg to define regularization losses. It leads to significantly improved synthetic shapes under novel poses.
@@
One way to address this issue is to learn generative models under explicit representations such as point clouds and meshes. Many approaches have shown promising results for point cloud generators for man-made shapes [193, 194, 195, 169, 196]. Although it is possible to fit a point cloud generator to shape collections with large geometric and topological variations, there are two limitations that prevent them from serving template models for fractured object reassembly. First, most approaches require the shapes to be consistently oriented, which only applies to a very limited number of categories. Generalization behavior under unaligned training instances is much worse. Second, the correspondences induced from the point cloud generator are not good enough across the synthetic shapes, making it difficult to properly define correspondences between the dynamic template and the input fragments. Recent results in mesh representation, such as deep marching tetrahedra [197], MeshGPT [198], and MeshAnything [199], also show impressive synthesis results. However, the two limitations of the point cloud representation remain under the mesh representation. An open question is how to address these two limitations.
6 Template-Based Reassembly
6.1 Complete object priors
Beyond purely feature-based approaches, using complete objects as templates can significantly boost the reassembly process. This is particularly valuable when working with well-documented objects like skulls, pottery, or architectural elements, where intact reference models often exist. Recent advances in machine learning have also made it possible to learn these geometric priors from collections of similar objects, offering more flexible template-based solutions.
@@ 6.1 Complete object priors
Methods using local-global alignment match fragmented pieces $x_{i}$ to the global prior $P$ and use this to obtain the initial or rough arrangement of the fragments. Their effectiveness depends on three factors: (1) the design of $P$, (2) the modeling/extraction of features from $\{x_{i}\}$ and $P$ for their matching, and (3) the alignment. The global prior $P$ could be a selected/computed explicit template or learned implicit parameters. Classical geometric reassembly methods often search for a good explicit shape and use it as a template or global prior, then design partial matching schemes to find local-global alignments. In contrast, as discussed in Sec. 6.3, recent learning-based methods tend to train neural networks to build implicit shape priors from datasets and then train/optimize the encoding of both local fragments and global template so that their latent codes can infer the matching and alignment of the fragments.
@@
In classical geometric methods, to support effective local-global alignment, the explicit shape template is often selected as an average or typical shape or a best-matched object retrieved from a database. Using a typical object as the prior shape is common in skull and cultural heritage reassembly tasks. For example, [204, 205, 200] used a standard human skull as $P$ to provide alignment guidance. [206] used a complete model from the museum’s collection as the prior to define the 3D support surface of the fragmented disk. When a larger collection of complete exemplar objects is available, retrieving a template object with the most similar geometry can provide a more adaptive global prior. For example, in the pottery reassembly field, researchers have built many digital databases of pottery. They support various types of local-global matching for the given fragments, such as drawing sketch [207], a portion of the object geometry [208], and curve descriptors [90, 209] to retrieve relevant potteries as priors $P$. In the archaeological and architectural reconstruction of a ruined monument task, to construct the most similar geometry as prior $P$, [51, 210] conducted a careful study of historical texts that mention the building under consideration. Given the selected shape prior $P$, as for the second factor, feature extraction computes keypoints and their descriptors on both fragments and the global prior, such as depth and axial symmetry descriptor [207], salient points [208], parabolic contour [90], spin images [204, 205], Signature of Histograms of OrienTations (SHOT) descriptor [200]. With the extracted keypoints and features, fragment transformations are computed to roughly assemble the input parts to the template. The researchers developed various alignment methods using keypoint correspondences, such as iterative closed points (ICP) [204], geometric distance [205], local affine-invariant 3D moments [90], Earth mover distance [208], and the RANSAC algorithm [200].
@@
Implicit prior for local-global alignment is learned from an annotated 3D dataset to solve the partial matching from the fragments to the global shape. The shape prior is encoded in a neural network, which is trained on a 3D segmentation dataset, and the weights are adjusted to learn the semantic relationship between the fragments and the complete objects. In the inference stage, given a set of parts $X$, trained networks are used as encoders to encode the parts into latent codes, which inherit the implicit shape prior to the possible complete object $P$. Chaudhuri et al. [211] trained a Bayesian network on a 3D segmentation dataset to encode semantic and geometric relationships among shape components. Then a probabilistic model is used to compose and align the input parts. Mo et al. [201] proposed a more comprehensive 3D object dataset, named PartNet, a large-scale dataset of 3D objects annotated with hierarchical and instance-level 3D part information. Following the PartNet dataset, Zhan et al. [212] and Narayan et al. [213] trained an encoder on it to encode the query parts to semantic latent codes, which inherit the shape prior to the complete object. They framed the 3D part assembly problem as a graph-learning problem. They took the latent codes as input and used graph neural networks to assemble the parts into complete objects. They demonstrated their work on 3D chairs, tables, and lamp reassembly tasks.
@@
Global evaluation as the complete object prior $P$ provides a regularization term to constrain the reassembled results. It has two variants: hand-crafted functions or learnable discriminators. Hand-crafted metrics are widely adopted in earlier reassembly tasks. In the image stitching task, an entropy-based metric is proposed to assess the quality of the stitched panoramic image [214]. In the document reassembly task, Optical Character Recognition (OCR) is used for word or sentence detection [215, 216]. These methods use pair-wise features to suggest candidate alignments of the fragments. They keep the candidate alignments when the evaluation results are promising, while they discard the alignments when the evaluation results are poor. More recent approaches leverage large-scale datasets to train neural networks that can evaluate assembly quality. These networks learn to distinguish between correct assemblies and incorrect ones, helping guide the reassembly process toward more plausible configurations. Li et al. [202] propose a GAN to solve the Jigsaw puzzles. They designed a network-based discriminator and an adversarial loss to classify the an image is reassembled or real. The adversarial loss pushes the GAN to assemble images from pieces that look natural. Then a discriminator was trained as adversarial loss to optimize the mating configuration that should look visually realistic like an object.
@@
Each approach offers distinct advantages for template-based reassembly. Local-global alignment is particularly effective when working with well-defined object categories that have representative complete objects or comprehensive reference databases. The success of these methods depends heavily on how effectively they can capture relationships between fragments and complete objects. Meanwhile, global evaluation approaches provide flexible guidance for assembly without requiring explicit templates, although the design of effective evaluation metrics remains an active research challenge. Future advances may come from combining these complementary approaches.
@@ 6.2.1 Symmetry Detection
Axial symmetry and reflectional symmetry are the two commonly studied symmetry types in geometry analysis and processing, especially in 3D reassembly. For a point cloud $P$, symmetry detection aims to find a transformation $s$ (either axial or reflectional) that minimizes:
$E(s)=\sum_{p\in P}\min_{q\in P}||s(p)-q||^{2}+\lambda R(s)$ (3)
where $s(p)$ is the transformed point under symmetry transform $s$, $q$ is its closest point in $P$, and $R(s)$ is a regularizer that ensures the transformation is a valid symmetry. For axial symmetry, $s$ represents rotation around an axis, while for reflectional symmetry, $s$ represents reflection across a plane.
As shown in Figure 14 (top), axially symmetric, also referred to as rotationally symmetric, indicates that if the complete object’s exterior surface is intersected by a plane perpendicular to its axis, the resulting cross section is a circle or closely resembles one. As a prominent geometrical property, rotational symmetry can serve as a feature to guide the 3D reassembly of fragmented relics, such as sherds. Research on symmetry axis estimation has evolved significantly over time. Early approaches focused on using differential geometric properties, including the sphere of curvature [220] and normal intersection [221] methods. Subsequently, enhancements of the estimated rotational axis have been proposed using M-estimators [222] or Bayesian approaches [223, 224]. These early methods, while foundational, often struggled with real-world challenges such as geometric deformation, missing pieces, and surface noise - common issues when working with fragmented objects. [225] suggested a workflow for axis estimation based on minimum thickness direction. This method does not rely on differential geometric properties, making it resilient to local defectiveness, noise, and outliers. Recent studies [226, 227, 217] have broadly adopted the random sample consensus algorithm (RANSAC) to find good initial estimates. [226] extended the Cao and Mumford’s method [220], to consider both the inner and outer surfaces. This technique introduces PotSAC, a two-stage axis estimator based on a variant of the RANSAC algorithm followed by a robust nonlinear least-square refinement, making it capable of handling noisy surface normals. [227] applied beam search to explore multitudes of registration possibilities, addressing false matches between sherds inflicted by indistinctive sharp fracture surfaces. [217] showed that the identification of the symmetry axis from surface normal lines may yield two meaningful local minima. A multiple solution RANSAC algorithm is proposed for initial estimates, followed by a coordinate descent algorithm to refine these estimates. The authors showed that their method ensures convergence, achieves a faster convergence rate, and provides state-of-the-art performance in axis estimation.
As illustrated in Figure 14 (bottom), reflection symmetry refers to the property that an object does not change upon undergoing reflection. The human skull typically exhibits reflection symmetry along the mid-sagittal plane [228]. This kind of symmetry is one of the main geometric priors used in skull reassembly and completion [205, 204].
@@ 6.2.2 Symmetry-guided reassembly
With an estimated symmetry axis, the consistency of the symmetry axis and the axis profile curve can be viewed as a constraint or optimization objective during reassembling. [229, 230] implemented a global constraint that all fragments share the same axis, greatly reducing computational costs and improving noise resistance. This method reduces the search space from 6 DOF (for two fragments) to 1 DOF (relative position of two pieces along the axis of symmetry). [147] exploits the results of an ongoing reassembly to predict the shape of the resulting object using symmetry-based expansion and then performs alignment of disjoint parts with the predicted shape to also assemble non-contacting parts. [227] relaxed the above constraint by encoding the information about the axis as a descriptor into each break line for efficient reassembly.
@@ 6.3 Learned Shape Spaces
Compared to approaches that rely on similar shapes or symmetric templates to guide fractured object reassembly, few methods have explored the use of a generic template shape space learned from the data to enhance this process. One notable exception is TAssembly [218], which learns a linear shape space specifically for bones. It formulates fractured object reassembly as an optimization problem, adjusting fragment poses and the underlying complete shape in the learned shape space to minimize the distance between the fragments and the dynamic template. However, this method is limited to bone shape spaces, where consistent mesh parameterization makes it relatively straightforward to learn descriptors for predicting dense correspondences. Another promising approach is Jigsaw++ [219], which employs a point cloud generation model with a rectified flow-based retargeting algorithm to infer the complete shape from a partially assembled input. Although this approach shows significant potential for reconstructing daily objects, scaling it to handle more diverse or even open-vocabulary input shapes requires larger models and more data. Significant challenges remain in developing effective shape priors for objects with diverse geometries and topologies. Additionally, current methods often lack mechanisms for establishing correspondences between reconstructed shapes and input fragments, making it difficult to enforce rigidity constraints between fragments.
@@ 7 Applications
We discuss four main applications areas for fractured object reassembly (Section 7.1 to Section 7.3) and the relevant application in shape modeling via part assembly.
@@ 7.1 Bone Reduction
3D bone reduction is essential for preoperative surgical planning [233, 234, 235] and intraoperative navigation [236, 237], including robotic surgery [238, 239](see Figure 16). The objective of bone reduction is to restore the anatomical alignment of fractured bones. Bone fragments have two unique characteristics that distinguish them from other fractured objects: First, human bones have strong shape priors due to their anatomical nature, including bilateral symmetry which allows the use of contralateral bones as templates (see Section 6). Second, bones have complex internal structures including cortical bone, cancellous tissue, and cartilage. When fractures occur in cancellous regions, especially in metaphyseal areas of long bones, the irregular fracture surfaces can make precise alignment challenging [218, 240].
Skull reconstruction, as a specialized case in bone reduction, presents unique characteristics and challenges. Unlike long bones, skull fragments are typically modeled as a collection of surfaces, where boundary loops serve as key features for matching [204, 205]. Feature descriptors computed on these boundary curves enable reliable fragment matching. The reconstruction process leverages the skull’s natural bilateral symmetry [228], particularly useful when fragments are missing from one side. For cases involving multiple feature curves on a single fragment, the matching problem can be formulated as the Largest Common Point-set (LCP) problem [200]. Statistical shape models trained on skull datasets provide additional guidance for cases with severe damage [218]. The precision requirements are particularly stringent due to the skull’s critical role in both surgical planning and forensic analysis.
The bone reduction workflow typically consists of four steps: CT bone segmentation, template matching, pairwise fragment matching, and multi-fragment matching.
CT bone segmentation endeavors to separate fractured bones from surrounding healthy tissue and divides individual fragments in CT images. A key challenge is distinguishing between fracture zones and intact zones. This can be done using automatic methods like watershed [233], probabilistic watershed [241], template-matching [240], or semi-automatic approaches like region growing [242, 243] and continuous max-flow [236]. Recent deep learning approaches include supervised contrastive learning [244], synergistic segmentation with contralateral bone templates [240] (see Figure 17), and dual-stream learning [245]. The outer surface of the intact zone [241, 240] serves as a template and can be extracted using the MarchingCube algorithm [48].
Template matching provides initial coarse alignment between bone fragments and a template. Most commonly, the healthy contralateral bone serves as a template [233, 240, 246, 241, 247, 248], though this approach has limitations due to anatomical variations [249] and may not always be available. As an alternative, parametric statistical shape models (SSM) [218, 234, 236] can be used. For complex cases involving joint dislocations, statistical pose models (SPM) [250] have been developed. A notable advancement by [236] combines both SSM and SPM in an iterative procedure that alternates between fragment alignment, SSM adaptation, and SPM adaptation.
@@
Pair-wise matching aligns two fragments by matching their fracture boundaries, involving fracture boundary identification followed by boundary matching. Early pioneering work [251] introduced a computational framework combining geometric analysis with medical imaging for pairwise bone fragment registration. After template matching brings fragments into rough alignment, fracture boundaries can be identified using Euclidean distance [233]. For fractures in the cortical bones, pronounced curvature helps to identify boundaries [246, 252, 253, 254, 241, 255]. However, fractures in trabecular bone often appear planar, requiring manual annotation [256, 246]. For cortical bone fractures, matching techniques similar to those used in the restoration of cultural artifacts [85, 242, 254] are employed, using criteria like tangential contact points and opposing surface normals. ICP is also commonly used [251, 246, 253, 255, 233]. For trabecular bone fractures or cases with missing fragments, template matching becomes essential.
Multi-fragment matching optimizes the poses of multiple fragments globally, using poses from template and pair-wise matching as input. The goal is to maximize region matching scores [200] through either exhaustive [242] or greedy search [257]. The greedy search method incorporates heuristic rules for matching and nex-bone selection, where various criteria, including template and pair-wise matching scores [200], fragment-to-set distance [252], fracture surface area [255], and fragment size [233]. To avoid local minima, beam search can select multiple fragments simultaneously [200]. Error accumulation is addressed through the ideal mates concept [252, 241, 257] or graph optimization [200]. Similar to stone fragment matching, bone reduction approaches often incorporate domain-specific knowledge.